# A Comprehensive Survey on Self-Supervised Learning for Recommendation

### v2

@@ 4.8.3. Generative Learning in MbeRec
Generative learning in MbeRec (Table 8) adopts the variational autoencoding paradigm to reconstruct behavior interactions, thereby providing informative auxiliary learning signals.
@@
  • Generative Learning Paradigm. Both VCGAE (Liu et al., 2023a) and BVAE (Rao et al., 2023) adopt the variational auto-encoding generative paradigm for modeling behavior heterogeneity. In VCGAE, a variation graph autoencoder is designed. It utilizes auxiliary behavior fusion to obtain the variance vector and a behavior transfer network to obtain the mean vector, which together form the distribution of latent factors. BVAE uses a behavior-aware semi-encoder to encode the variance vector based on users’ historical interactions. It learns the mean vector through a global feature filtering network, resulting in behavior-aware latent representations of users.
@@
  • Generation Target. In the VAE paradigm, the latent representations obtained are fed into the decoder for generation. In VCGAE, the generation target is the edge in the user-item graph of the target behavior (e.g., purchase). Besides, BVAE focuses on generating the user interaction vector across different behaviors.
@@ 4.9.2. Contrastive Learning in MmoRec
  • View Creation. In data-based view creation methods, CMI (Li et al., 2022a) conducts data-level item sequence augmentation and encodes user interest embeddings to generate contrastive views. MMGCL (Yi et al., 2022) employs modality masking and edge dropout techniques to create augmented data views for each node in the multi-modal graph. Additionally, SLMRec (Tao et al., 2022b) utilizes data-level augmentation techniques such as feature dropout and masking to obtain two views for each node. In contrast, model-based methods like Liu et al. (Liu et al., 2022b) leverage a text/image encoder to encode two views for each item, which are then used for contrastive learning. Furthermore, MICRO (Zhang et al., 2022d) mines the latent graph structure of items and utilizes graph neural networks to encode modality-specific views and generate a fused view for each item. Meanwhile, MMSSL (Wei et al., 2023a) employs adversarial learning to obtain a modality-aware user-item graph and encodes modality-fused user embeddings as one view and another user’s view based on collaborative information. For feature-based view creation, BM3 (Zhou et al., 2023b) encodes modality embeddings, applies feature-level embedding dropout for augmented views, and includes an ID-based view from ID embeddings.
@@
  • Pair Sampling. All the methods follow natural sampling. In (Liu et al., 2022b), positive pairs are formed by two views from different modalities of the same item, and negative pairs consist of views from different items. CMI creates positive pairs from two augmented interest-based views for the same user. MMGCL forms positive pairs from randomly augmented views and challenging negatives obtained by replacing modality data. SLMRec forms positive pairs with two data-augmented views of each node and ID-modality and other modality representations. MICRO maximizes agreement between item representations under individual modalities and fused multi-modal representations, forming positive pairs with modality-view and fused view of the same item and negative pairs with views from other items. BM3 aligns inter-modality by pairing augmented modality-views with the ID-based view and intra-modality by pairing modality-views of the same item. MMSSL forms positive pairs with the fused modality-view and collaborative-aware view of the same user, and negative pairs with views from other users.
@@
  • Contrastive Objective. Most of the CL-based works utilize either InfoNCE-based (Li et al., 2022a; Yi et al., 2022; Tao et al., 2022b; Zhang et al., 2022d; Wei et al., 2023a) or JS-based (Liu et al., 2022b) loss functions as the loss function. Besides, BM3 directly aligns the views with cosine similarity.
4.9.3. Generative Learning in MmoRec
@@
Generative learning, as recently shown by MVGAE (Yi and Chen, 2021) (Table 8), employs the graph VAE to transform modality data into a latent space and utilizes generation tasks to guide the learning process.
@@
  • Generative Learning Paradigm. MVGAE utilizes a multi-modal variational graph auto-encoder to assign mean and variance vectors to each node in the user-item graph and employs the product-of-experts approach to fuse modality-specific latent representations for generation.
@@
  • Generation Target. In MVGAE, the latent representations obtained from encoder are utilized to generate (reconstruct) the interaction edges in the user-item graph using an inner product decoder and BPR loss.
@@ 4.9.4. Adversarial Learning in MmoRec
  • Adversarial Learning Paradigm. MMSSL generates a modality-aware user-item graph structure using a neural generator with item multi-modal features. Then, the user-specific interaction vector undergoes adversarial learning in a discriminator. The differentiable pipeline enables multi-modal adversarial user-item relation learning.
  • Discrimination Target. The user-item graph structure learned in MMSSL is split into user-specific interaction vectors, which are used for discrimination. The objective is to make these interaction vectors similar to real user interaction records in order to deceive the discriminator.
@@ 5.1. Towards Foundation Recommender Models
Foundation model (Bommasani et al., 2021), trained on massive datasets and exhibiting remarkable generalization capabilities to handle a wide range of downstream tasks, has garnered significant attention from researchers across various domains (Radford et al., 2021; Bubeck et al., 2023; Xia et al., 2024; Tang et al., 2023). In recommender systems, current models using SSL techniques have shown significant performance, but they are limited by evaluation settings focused on a single dataset. Recent research in CV and NLP has improved models’ generalization abilities through self-supervised learning using contrastive learning (Radford et al., 2021) or generative learning (Touvron et al., 2023) on large amounts of data. A promising future direction is to design foundation recommender models that leverage SSL to learn user-item interaction patterns from massive data and achieve zero-shot cross-data reasoning and recommendation.
@@ 5.2. Unleashing the Potential of Denoised Diffusion Models
Denoised diffusion models have demonstrated remarkable proficiency in generating diverse types of data, including images, text, and even structured data. This new and popular generative learning paradigm has been effectively applied in various domains, as evidenced by recent studies such as (Rombach et al., 2022; Gong et al., 2022), and (Vignac et al., 2022). Recent studies in recommendation with generative self-supervised learning have started leveraging diffusion models either to generate augmented data (Liu et al., 2023b) or as the backbone model for inference (Wang et al., 2023d). We believe that the impressive generation capabilities demonstrated by diffusion models will bring new insights and fascinating models to the generative learning-enhanced recommendation in the near future, making it a promising research line.
@@ 5.3. Self-Supervised Integration of Large Language Models
Large language models (LLMs) have gained significant attention due to their exceptional performance in various domains (Zhao et al., 2023; Tang et al., 2024; Liu et al., 2024a). In recommender systems, LLMs can generate user/item profiles (Ren et al., 2024a; Wei et al., 2024a), explain user-item interactions (Li et al., 2023f), and serve as the backbone for recommendation with instruction tuning (Zhang et al., 2023a). The integration of large language models has brought about an unprecedented abundance of diverse, rich, and high-quality textual modality data. However, the challenge of effectively harnessing this textual capability remains an open research question. Recent works have employed self-supervised learning techniques, such as contrastive learning (Ren et al., 2024a) and mask-reconstruction (Wei et al., 2024a), to align the knowledge of LLMs with recommenders. As we look to the future, self-supervised learning techniques are poised to play a crucial role in enhancing recommenders with the incorporation of LLMs.
@@ 5.4. Self-Supervised Learning for Dynamic Recommendation Adaptation
5.5. Theoratical Foundation of various SSL paradigm
While self-supervised learning has improved recommender systems, a comprehensive theoretical foundation for various paradigms is lacking. Some methods have provided theoretical explanations for contrastive learning (Wu et al., 2021b; Ren et al., 2023b) and generative masked autoencoding in graphs (Li et al., 2022c). However, recent paradigms like denoised diffusion still need theoretical explanations to demonstrate their benefits in recommendation. This understanding would help uncover the underlying principles and mechanisms driving self-supervised learning in recommendation systems and offer insights into the generalization and robustness properties of these algorithms.

### v3

@@ 4.8.3. Reconstructive Learning in MbeRec
Reconstructive learning in MbeRec (Table 8) adopts the variational autoencoding paradigm to reconstruct multi-behavior interactions, thereby providing informative auxiliary learning signals for behavior-aware recommendation.
@@
  • Reconstructive Learning Paradigm in MbeRec leverages variational autoencoding mechanisms to model behavior heterogeneity and capture complex interaction patterns. Both VCGAE (Liu et al., 2023b) and BVAE (Rao et al., 2023) adopt the variational auto-encoding generative paradigm for modeling behavior heterogeneity. VCGAE designs a variational graph autoencoder that utilizes auxiliary behavior fusion to obtain the variance vector and employs a behavior transfer network to derive the mean vector, jointly forming the distribution of latent factors. BVAE uses a behavior-aware semi-encoder to encode the variance vector based on users’ historical interactions and learns the mean vector through a global feature filtering network, resulting in behavior-aware latent user representations.
@@
  • Reconstruction Target here focuses on reconstructing behavior-specific patterns to enhance representation learning. In the VAE paradigm, the obtained latent representations are fed into the decoder for generation tasks. VCGAE targets the reconstruction of edges in the graph for specific target behaviors (e.g., purchase), while BVAE focuses on generating comprehensive user interaction vectors across different behavioral modalities.
4.8.4. In-depth Analysis of SSL-based Multi-behavior Recommendation
In multi-behavior recommendation, CL-based methods excel at modeling behavior heterogeneity by creating behavior-specific views and leveraging multi-behavior interactions (Yang et al., 2021a; Wu et al., 2022b; Xu et al., 2023). These methods effectively capture inter- and intra-behavior information, enabling fine-grained representation learning. Meanwhile, reconstructive methods (Liu et al., 2023b; Rao et al., 2023) offer an alternative by focusing on behavior reconstruction to uncover latent patterns. While CL-based methods dominate for their flexibility and multi-view learning, integrating generative paradigms could further improve the robustness of multi-behavior recommendation models in the future.
@@ 4.9.2. Contrastive Learning in MmoRec
  • View Creation in MmoRec methods constructs meaningful representations from multi-modal item features through data-based, model-based, and feature-based strategies. In data-based view creation methods, CMI (Li et al., 2022a) conducts data-level item sequence augmentation and encodes user interest embeddings to generate contrastive views. MMGCL (Yi et al., 2022) employs modality masking and edge dropout for augmented graph views. Additionally, SLMRec (Tao et al., 2022b) utilizes data-level augmentation techniques such as feature dropout and masking to obtain two views for each node. In contrast, model-based methods like Liu et al. (Liu et al., 2022c) leverage a text/image encoder to encode two views for each item, which are then used for contrastive learning. Furthermore, MICRO (Zhang et al., 2022a) mines the latent graph structure of items and utilizes graph neural networks to encode modality-specific views and generate a fused view for each item. Meanwhile, MMSSL (Wei et al., 2023a) employs adversarial learning to obtain a modality-aware user-item graph and encodes modality-fused user embeddings as one view and another user’s view based on collaborative information. For feature-based view creation, BM3 (Zhou et al., 2023b) encodes modality embeddings, applies feature-level embedding dropout for augmented views, and includes an ID-based view from ID embeddings.
@@
  • Pair Sampling in MmoRec methods constructs positive and negative pairs from multi-modal representations and augmented views. All the methods follow natural sampling. In (Liu et al., 2022c), positive pairs are formed by two views from different modalities of the same item, and negative pairs consist of views from different items. CMI creates positive pairs from two augmented interest-based views for the same user. MMGCL forms positive pairs from randomly augmented views and challenging negatives obtained by replacing modality data. SLMRec forms positive pairs with two data-augmented views of each node and ID-modality and other modality representations. MICRO maximizes agreement between item representations under individual modalities and fused multi-modal representations, forming positive pairs with modality-view and fused view of the same item and negative pairs with views from other items. BM3 aligns inter-modality by pairing augmented modality-views with the ID-based view and intra-modality by pairing modality-views of the same item. MMSSL forms positive pairs with the fused modality-view and collaborative-aware view of the same user, and negative pairs with views from other users.
@@
  • Contrastive Objective in MmoRec systems utilizes InfoNCE-based (Li et al., 2022a; Yi et al., 2022; Tao et al., 2022b; Zhang et al., 2022a; Wei et al., 2023a) or JS-based (Liu et al., 2022c) loss functions, while BM3 directly aligns views with cosine similarity.
4.9.3. Reconstructive Learning in MmoRec
@@
Reconstructive learning, as recently shown by MVGAE (Yi et al., 2021) (Table 8), employs the graph VAE to transform modality data into a latent space and utilizes generation tasks to guide the learning.
@@
  • Reconstructive Learning Paradigm in MmoRec integrates multi-modal information through variational autoencoding, where MVGAE assigns mean and variance vectors to each user-item graph node and employs product-of-experts to fuse modality-specific latent representations.
@@
  • Reconstruction Target here reconstructs graph structural information, where MVGAE uses encoder-derived latent representations to reconstruct user-item interaction edges via inner product decoder and BPR loss.
@@ 4.9.4. Adversarial Learning in MmoRec
  • Adversarial Learning Paradigm in MmoRec generates user-item structures through multi-modal adversarial mechanisms. MMSSL generates a modality-aware user-item graph structure using a neural generator with item multi-modal features. Then, the user-specific interaction vector undergoes adversarial learning in a discriminator.
  • Discrimination Target here distinguishes real and generated user-item interactions, where the user-item graph structure learned in MMSSL is split into user-specific interaction vectors, which are used for discrimination.
4.9.5. In-depth Analysis of SSL-based Multi-modal Recommendation
Three SSL techniques each have distinct attributes. Contrastive learning methods (Li et al., 2022a; Zhou et al., 2023b) excel at utilizing diverse views of multi-modal data to enhance representation learning; however, they may face challenges in creating high-quality multi-modal views for optimization. Reconstructive methods (Yi et al., 2021) and adversarial learning approaches (Wei et al., 2023a) both provide robust capabilities: the former captures complex interdependencies among modalities, while the latter offers a dynamic framework for modeling intricate user-item interactions, promoting more nuanced embeddings. Both approaches require the appropriate design of reconstruction objectives and adversarial processes to effectively harness multi-modal data.
@@ 5.1. Towards Foundation Recommender Models
Foundation model (Bommasani et al., 2021), trained on massive datasets and exhibiting remarkable generalization capabilities to handle a wide range of downstream tasks, has garnered significant attention from researchers across various domains (Radford et al., 2021; Bubeck et al., 2023; Xia et al., 2024; Tang et al., 2024a). In recommender systems, current models using SSL techniques have shown significant performance, but they are limited by evaluation settings focused on a single dataset. Recent research in CV and NLP has improved models’ generalization abilities through self-supervised learning using contrastive learning (Radford et al., 2021) or generative learning (Touvron et al., 2023) on large amounts of data. A promising future direction is to design foundation recommender models that leverage SSL to learn user-item interaction patterns from massive data and achieve zero-shot cross-data reasoning and recommendation.
@@ 5.2. Unleashing the Potential of Denoised Diffusion Models
Denoised diffusion models have demonstrated remarkable proficiency in generating diverse types of data, including images, text, and even structured data. This new and popular generative learning paradigm has been effectively applied in various domains, as evidenced by recent studies such as (Rombach et al., 2022; Gong et al., 2022), and (Vignac et al., 2022). Recent studies in recommendation with generative self-supervised learning have started leveraging diffusion models either to generate augmented data (Liu et al., 2023a) or as the backbone model for inference (Wang et al., 2023e). We believe that the impressive generation capabilities demonstrated by diffusion models will bring new insights and fascinating models to the generative learning-enhanced recommendation in the near future, making it a promising research line.
@@ 5.3. Self-Supervised Integration of Large Language Models
Large language models (LLMs) have gained significant attention due to their exceptional performance in various domains (Zhao et al., 2023; Tang et al., 2024b; Liu et al., 2024a). In recommender systems, LLMs can generate user/item profiles (Ren et al., 2024a; Wei et al., 2024a), explain user-item interactions (Li et al., 2023e), and serve as the backbone for recommendation with instruction tuning (Zhang et al., 2024b). The integration of large language models has brought about an unprecedented abundance of diverse, rich, and high-quality textual modality data. However, the challenge of effectively harnessing this textual capability remains an open research question. Recent works have employed self-supervised learning techniques, such as contrastive learning (Ren et al., 2024a) and mask-reconstruction (Wei et al., 2024a), to align the knowledge of LLMs with recommenders. As we look to the future, self-supervised learning techniques are poised to play a crucial role in enhancing recommenders with the incorporation of LLMs.
@@ 5.4. Self-Supervised Learning for Dynamic Recommendation Adaptation
5.5. Theoretical Foundation of various SSL paradigm
While self-supervised learning has improved recommender systems, a comprehensive theoretical foundation for various paradigms is lacking. Some methods have provided theoretical explanations for contrastive learning (Wu et al., 2021a; Ren et al., 2023a) and generative masked autoencoding in graphs (Li et al., 2023c). However, recent paradigms like denoised diffusion still need theoretical explanations to demonstrate their benefits in recommendation. This understanding would help uncover the underlying principles and mechanisms driving self-supervised learning in recommendation systems and offer insights into the generalization and robustness properties of these algorithms.
@@ 6. Conclusion
This survey systematically reviews self-supervised learning (SSL) frameworks for recommender systems, addressing data sparsity challenges and enhancing predictive accuracy. Through analysis of hundreds of studies, we establish a comprehensive taxonomy categorizing SSL paradigms including contrastive, generative, and adversarial learning. Our findings highlight SSL’s transformative role in improving personalization and user experience in recommendation systems, serving as a foundational resource for researchers developing state-of-the-art recommender systems.
Acknowledgments
This work is supported by the National Natural Science Foundation of China under Grants 624B2122.